The length of a rectangle is 1cm longer than its width. If the diaganol of rectangle is 4cm, what are the dimensions (length and width) of the rectangle?
Thank you

Answer by checkley71(8403)

using the pathagorian theorum a^2+b^2=c^2 a=x, b=(x+1), c=4
x^2+(x+1)^2=4^2
x^2+x^2+2x+1=16
2x^2+2x-15=0 using the quadratic equation x=(-b+-sqrt[b^2-4ac])/2a
x=(-2+-sqrt[2^2-4*2*-15])/2*2
x=(-2+-sqrt[4+120])/4
x=(-2+-sqrt124)/4
x=(-2+-11.1355)/4
x=(-2+11.1355)/4
x=9.1355/4
x=2.2839 cm for the width.
length is 2.2839+1=3.2839
proof
2.2839^2+3.2839^2=4^2
5.216+10.784=16
16=16
#########################################
HAPPY NEW YEAR.